Evaluate LaTeX: \int_cF.dr\:\: where LaTeX: F\left(x,y,z\right)=xzi-yzkF(x,y,z)=xzi−yzk and c is the line segment from (3,0,1) to (-1,2,0)


1
Expert's answer
2021-05-04T13:04:50-0400

Given, "F(x,y,z)=xz\bold{i}-yz\bold{k}" and line segment c from (3,0,1) to (-1,2,0).

Then,

"r(t)=(1-t)<3,0,1>+t<-1,2,0>=<3-4t,2t,1-t>\newline
\text{Differentiate with respect to t,}
\frac{dr}{dt}=<-4,2,-1>\newline
dr=(-4i+2j-k)dt\newline
\text{Now, }F(r(t))=(3-4t)(1-t)i-2t(1-t)k\newline
\hspace{2.1cm}=(3t^2-7t+3)i+(2t^2-2t)k\newline
\text{Therefore,}\newline
\intop F.dr=[(3t^2-7t+3)i+(2t^2-2t)k].[−4i+2j−k]dt\newline
\hspace{1.1cm}=\int^1_{0}-4(3t^2-7t+3)-(2t^2-2t)dt\newline
\hspace{1.1cm}=\int^1_{0}-14t^2+30t-12dt\newline
\hspace{1.1cm}=[-\frac{14}{3}t^3+15t^2-12t]^1_{0}\newline
\hspace{1.1cm}=\frac{-5}{3}"

Thus,the requrired answer is 5/3.